OpenStudy (anonymous):

Can I pleaseeeeee get some help with some math? I'll post a link to the first question below.

OpenStudy (anonymous):

http://gyazo.com/00293db4925b1c41cb1305f9845eb742

OpenStudy (anonymous):

I don't know how, sorry.

OpenStudy (anonymous):

The parallel lines make the two triangles similar. The sides are split proportionally. You could say \[\frac{9}{4x-2}=\frac{12}{3x+2}\] and use the properties of proportions to solve it. (namely, cross products property)

OpenStudy (anonymous):

Okay..I think I got that, how would I use the properties of proportions?? Like, how would I solve it?

OpenStudy (anonymous):

The cross products property says that for the proportion \[\frac{a}{b}=\frac{c}{d}\] the "cross-products" are equal. That is, a*d (the first cross product) = b*c (the other cross product).This can be done equally as easily with elementary algebra, but this property simplifies it. For a/b=c/d, ad=bc. Plug in the a, b, c, and d values from your homework and you can solve it like any other equation.

OpenStudy (anonymous):

Uh. I guess I got it, thanks.

OpenStudy (anonymous):

Maybe sometime you can help me do it. Cause I still can't, even though I slightly understand what you mean. I just. Can't. Im horrible at math. It looks like soup to me.

OpenStudy (anonymous):

I have got to go to bed. I'll reply to anything in the morning.

OpenStudy (whpalmer4):

\[9(3x+2) = 12(4x-2)\]use distributive property: \[9*3x + 9*2 = 12*4x - 12*2\]\[27x+18=48x-24\]

OpenStudy (anonymous):

Again, thank you very much!!! Both of you.

OpenStudy (whpalmer4):

Now subtract 27x from both sides, and add 24 to both sides